OpenStudy (anonymous):

Find the values of x in the following equation: |-5x + -2| = 12

OpenStudy (anonymous):

-5x + -2 = 12 -5x = 14 x = -14/5 or -(-5x - 2) = 12 5x + 2 = 12 5x = 10 x = 2 two roots: 2 , -14/5

OpenStudy (anonymous):

the | 's means the modulus of the function, meaning if the function inside the |'s is negative the output is positive

OpenStudy (anonymous):

so if the function inside is positive or negative the value is

OpenStudy (anonymous):

ok guy?

OpenStudy (anonymous):

k